The frequency of an allele p = 0.1. If the viabilities of the three genotypes would be w11 = 0.9, w12 = 0.95 and w22 = 1. what would be the frequency of the allele in the next generation?
Answer #1

We have w11 =:0.9 = p2

w12=0.95 = 2pq

w22 = 1 = q2

The allele frequency in the next generation will be

p2 + pq = 0.9 + 0.95/2 = 0.9 + 0.475

= 1.375 .